OpenStudy (anonymous):

Which of the linear systems has infinitely many solution? A) 15x+y=20 6x-2y=8 B) 15x+5y=20 6x-2y=-8 C) 15x+5y=-20 6x-2y=8 D) 15x-5y=20 6x-2y=8

OpenStudy (anonymous):

help me pls guys?...

OpenStudy (btaylor):

If a system of equations has infinitely many solutions, then the two lines coincide. So, if you were to use linear combination (as the format of this problem suggests), it would cancel out to get 0=0 or something like that. I'd figure out how to combine each set of equations and see which one turn out to be 0=0.

OpenStudy (anonymous):

B

OpenStudy (btaylor):

nope.

OpenStudy (anonymous):

its wrong

OpenStudy (anonymous):

A because that's the right

OpenStudy (btaylor):

If you combined A, you would get: 15x + y = 20 -15x + 5 y = -20 -6y=0 You want it to get 0=0.

OpenStudy (anonymous):

so i'm right

OpenStudy (anonymous):

A system of 2 linear equations has an infinity of solutions when the 2 line graphs are coincident (or parallel), meaning they have the same slope. Look at system D. 5y = 15x - 20 --> y = 3x - 4 2y = 6x - 8 --> y = 3x - 4 The 2 lines are coincident, there is an infinity of solutions.